# Learning when to rank: Estimation of partial rankings from sparse, noisy comparisons

### v2

@@ II.4 Optimization
3:$L_{\text{current}}\leftarrow\text{ComputeLoss}(r,\sigma)$
4:$L_{\text{best}}\leftarrow L_{\text{current}}$
5:$r_{\text{best}}\leftarrow r$, $\sigma_{\text{best}}\leftarrow\sigma$
@@
We propose a fast nonparametric agglomerative algorithm based on an alternating optimization strategy to approximate the MAP estimates $\bm{\hat{r}}$, $\bm{\hat{\sigma}}$ for the node rankings and strengths. Starting from an initial condition in which each node belongs to its own group, so that $R=N$, the algorithm alternates between updating the unique strengths $\bm{\sigma}$ and rankings $\bm{r}$ so as to minimize the negative log-posterior $\mathcal{L}$ (Eq. II.3).
@@
Eq. 28 defines a set of $R$ self-consistent equations for the $\sigma_{r}$’s which can be iteratively solved until convergence. If $E(\bm{r})$ is the number of non-zero entries in the matrix $\bm{\omega}$ when the ranks are equal to $\bm{r}$, then the complexity of this algorithm is $O(E(\bm{r}))$. These iterative updates for the player’s strengths are of the same form as those introduced in Eq. 12 and benefit from the same convergence guarantees.
@@
which is the standard iterative update of Eq. 28 and can be solved efficiently with a computational complexity of $O(E(\bm{r})/R)$, where $E(\bm{r})$ is now the number of non-zero entries in the matrix $\bm{\omega}$ for the current ranks $\bm{r}$. When applied across all $R-1$ possible adjacent rank pairs $(r,r^{\prime})$, the total complexity for this update and merge process is then $O(E(\bm{r}))$.
@@
Fig. 1 shows the results of applying both the BT model and our Partial Rankings (PR) algorithm (with the BT likelihood of Eq. 3) to a small dataset of dominance interactions in a pack of wolves van198711 . We notice that our model offers a more concise interpretation of the data by grouping individuals with similar BT scores into the same rankings.
@@ III.1 Synthetic Match Datasets
Figure 2: Partial rankings in synthetic datasets. (a) Heatmap of the number of rankings $R$ inferred by the partial rankings model, as a function of the parameters $(\sigma,\langle k\rangle)$ of the synthetic model in Sec. III.1. (b) Heatmap of the log posterior odds ratio (Eq. 37) between the BT and the partial rankings model across the simulations. Positive values indicate a preference for the partial rankings model, and negative values a preference for the BT model. (c) The top panel displays the average number of rankings inferred by the best-performing model (blue) and our partial rankings model (red) as a function of the average degree $\langle k\rangle$ of the network of matches for the case in which no planted partial rankings are present. The bottom panel displays the log posterior odds ratio (normalized per node) between the two models. Negative values of this difference indicate a preference for the BT model and positive values a preference for the partial rankings model. All results were obtained by averaging over $20$ different simulations from the synthetic network model of Sec. III.1, and error bars indicate $2$ standard errors in the mean. Except near the transition point, error bars are smaller than the marker size.
@@
    For a given value of $\sigma$, assign three scores $[-\sigma,0,\sigma]$ (representing player strengths $[e^{-\sigma},1,e^{\sigma}]$) to define the planted partial rankings.
@@
By varying $M$ and $\sigma$, we generate networks of matches with different levels of sparsity and strength separation, allowing us to assess the performance of our algorithm across a wide range of scenarios. Results, shown in Fig. 2, depict the algorithm’s ability to recover the planted rankings as a function of $\sigma$ and the average degree of the network $\langle k\rangle=M/N$. For each $(\sigma,\langle k\rangle)$ pair, $20$ different networks were generated, and the results were then averaged. The ranges of values for $\langle k\rangle$ and $\sigma$ were chosen so as to capture all the relevant behaviors of the model.
As expected, we observe that for low values of $\langle k\rangle$, there is not enough evidence to recover the planted rankings regardless of the strength separation between players. In such sparse conditions, the algorithm assigns all players to the same rank, as shown in Fig. 2(a). This behavior contrasts sharply with the BT model, which still produces a complete ranking of the nodes even in these highly sparse conditions. As the number of matches increases and the networks become denser, the model gains sufficient statistical evidence to distinguish between the rankings and eventually recovers the planted structure. Notably, the number of matches required to correctly recover the rankings decreases rapidly with greater strength separation between players. In these cases, stronger players consistently outperform weaker ones, creating clearer signals that enable the ranking structure to be resolved with less data.
@@
as a function of $\langle k\rangle$ and $\sigma$. We observe that, when the separation in player strengths is small, our partial rankings model outperforms BT in sparse networks. In these cases, the PR model effectively leverages its ability to group players into shared ranks, avoiding overfitting. However, as the networks become denser, while our model accurately recovers the planted rankings even in this low-separation regime, it does not offer a more parsimonious description of the data (in terms of the posterior odds ratio of Eq. 37) compared to the standard BT model which, in contrast, infers a fully ranked structure for the nodes with $R=N=50$. This suggests a noisy regime where the evidence supporting partial rankings is relatively weak. The BT likelihood provides a more flexible fit for fine-gained variations in the data in these dense regimes (effectively overfitting the data), even if its prior poorly reflects the underlying data.
@@
In Fig. 2(c), we analyze the performance of both models in the absence of partial rankings, using networks of $N=150$ players, each assigned a unique score sampled from a logistic distribution with mean zero and scale one—a setting in which the BT model is expected to accurately capture the data. Averaged over 20 network realizations, the results show that for small values of $\langle k\rangle$, our partial rankings (PR) model offers a more parsimonious fit, as the limited number of interactions prevents reliable inference of a complete ranking. As $\langle k\rangle$ increases, however, we observe a sharp transition in which the BT model quickly surpasses the PR model in explanatory power. We also observe that the PR model continues to infer more rankings as $\langle k\rangle$ grows, albeit very slowly—indicating that, in principle, it is capable of recovering the full ranking given sufficient data.
@@
Interestingly, as the inferred number of groups increases with $\langle k\rangle$, the log posterior odds ratio decreases. This is somewhat counterintuitive, as one might expect the posterior odds ratio to improve as the number of rankings inferred by the PR model approaches the true number of rankings. However, as mentioned in Sec. II, the PR model tends to heavily penalize the presence of a large number of rankings so that the net effect of increasing the number of rankings is a reduction in the posterior odds ratio, with the PR model becoming a less efficient encoding for the rankings as the number of inferred groups grows. In practice, the best practice is to evaluate both an original ranking method and our corresponding partial ranking method to determine which provides a more parsimonious fit to the data in terms of log-posterior odds.
@@
To assess the goodness of the recovered partitions, in Appendix C we compute the Kendall rank correlation coefficient ($\tau_{B}$), adjusted for ties, between the ground truth rankings and the rankings inferred by the BT and PR models. This comparison is performed both when partial rankings are present and when they are not. We observe that when partial rankings are present in the network, the PR model can quickly attain perfect recovery as the networks become denser. On the other hand, while the BT model still achieves high $\tau_{B}$ scores, it never achieves perfect ordinal association and always underperforms relative to the PR model. However, when no partial rankings are present, and each player is assigned a unique score, the roles are reversed, and BT always displays a higher ordinal association with the ground truth rankings.
@@
We also compare our model with the partial rankings derived by applying 1D mean shift (MS) clustering comaniciu2002mean to the rankings inferred by the BT model, in Appendix D. We observe that our algorithm can always recover the correct number of partial rankings when these are present, while MS fails to do so. More importantly, our algorithm can adjust the number of inferred clusters as more data becomes available, even in cases when no partial rankings exist, resulting in a more accurate description of the rankings. In contrast, mean shift clustering consistently infers roughly the same number of partial rankings regardless of the evidence provided by the data.
@@ III.2 Real Match Datasets
Figure 3: Partial rankings in real networks of pairwise comparisons. (a) (top) Rescaled number of ranks per node inferred by the best-performing model and (bottom) log posterior odds ratio per node, both as a function of the average degree for all real-world networks considered in the study (see Table 1). The colours of the points indicate the different categories to which the datasets belong, while the shape of the markers indicates which model (PR or BT) emerged as the best-performing method (in terms of the posterior odds ratio). Positive values of the log posterior odds ratio indicate evidence in favour of the PR model; negative values indicate evidence in favour of the BT model. (b) Effective number of ranks, $R^{*}$ (Eq. 38), as a function of the number of unique ranks $R$ inferred by the PR algorithm. The black dashed line represents the line $R^{*}=R$. (c) Number of rankings inferred via the partial rankings algorithm as a function of the number of rankings inferred via mean shift clustering. Point colours indicate the $\tau_{B}$ score between the two inferred rankings. The black dashed line represents the line $R_{PR}=R_{MS}$. (d) Number of nodes of each network as a function of the number of rankings $R$ inferred by the partial rankings algorithm. Point colours indicate the $\tau_{B}$ score between the rankings inferred via the PR algorithm and those inferred by the BT model. The dashed black line indicates the line $R=N$.
Data set $N$ $M$ $\langle k\rangle$ $R$ $R^{*}$ $\Delta\mathcal{L}$ Category Description Ref.
Soccer 2204 7438 3.4 1 1 1469.3 Sports/games Men’s international association football matches 2010 - 2019 soccer
Friends 774 2799 3.6 2 2.0 450.1 Human High-school friend nominations high_school_friendships
Tennis 1272 29397 23.1 6 1.7 404.9 Sports/games Association of Tennis Professionals matches 2010–2019 tennis
Chess 917 7007 7.6 1 1 357.5 Sports/games Online chess games on lichess.com in 2016 chess
@@
Little Rock food web 183 2494 13.6 5 3.2 83.7 Animal Food web among the species in Little Rock Lake in Wisconsin martinez1991artifacts
@@
CS departments 205 34388 21.4 5 3.6 33.4 Human PhD graduates of one department hired as faculty in another clauset2015systematic
College football 115 593 5.2 1 1 26.8 Sports/games NCAA College Football matches 2013-2023 collegefootball
@@
Dutch school friends 26 234 9.0 2 1.3 9.2 Human Friendships at secondary school in The Netherlands, 2003-2004 snijders2010introduction
@@
Video games 125 1951 15.6 3 2.2 4.7 Sports/games Super Smash Bros Melee tournament matches in 2022 videogames
@@
History departments 144 4112 28.6 6 3.8 -3.2 Human PhD graduates of one department hired as faculty in another clauset2015systematic
@@
Hyenas 29 1913 66.0 9 7.9 -7.6 Animal Dominance interactions among hyenas in captivity strauss2019social
@@
Sparrows 26 1238 47.6 8 7.3 -15.4 Animal Attacks and avoidances among sparrows in captivity watt1986relationship
@@
Baboons 53 4464 84.2 13 10.3 -16.3 Animal Dominance interactions among baboons in captivity franz2015self
@@
Dogs 27 1143 42.3 6 5.3 -20.3 Animal Aggressive behaviors in a group of domestic dogs silk2019elevated
Wolf 15 10382 692.1 10 8.6 -24.0 Animal Dominance interaction among wolves in captivity van198711
@@
Mice 30 1230 41.0 5 4.2 -26.8 Animal Dominance interactions among mice in captivity williamson2016mouse
@@
Business departments 112 7856 70.1 9 7.3 -35.2 Human PhD graduates of one department hired as faculty in another clauset2015systematic
@@
Vervet monkeys 41 2980 72.7 8 6.8 -42.7 Animal Dominance interactions among a group of wild vervet monkeys vilette2020comparing
Scrabble 587 23477 40.0 2 2.0 -86.0 Sports/games Scrabble tournament matches 2004–2008 scrabble
Basketball 240 10002 41.7 2 2.0 -90.7 Sports/games National Basketball Association games 2015–2022 basketball
Soccer (aggregated) 288 7438 25.8 6 4.8 -131.1 Sports/games Soccer dataset above aggregated across 2010 - 2019 soccer
UN migrations 232 11228 48.4 194 183.2 -1036.5 Human Migrations between countries 1990 - 2015 unmigrations
@@
Table 1: Datasets analysed using the PR and BT algorithms, in order of decreasing log posterior odds ratio in Eq. 37 (i.e. increasingly in favor of the original BT model). $N$, $M$, and $\langle k\rangle$ indicate the number of nodes, edges, and average degree of each network, respectively. $R$ and $R^{*}$ indicate the number of partial rankings inferred via the PR algorithm and the effective number of partial rankings (Eq. 38). $\mathcal{L}$ indicates the logarithm of the posterior odds ratio between the PR and BT models (Eq. 37).
@@
In this section, we focus on how our PR algorithm may be used to derive meaningful patterns in real-world data. We consider the datasets compiled in jerdee2024luck and available at jerdeegit . In addition, we also consider a dataset of social hierarchies among a pack of wolves van198711 ; newman2023efficient , and other sets of directed networks available at netzschleuder . The datasets are summarized in Table 1.
@@
We apply our PR algorithm and the BT model to all match lists in the dataset and compare the results. In Fig. 3(a), we show the number of inferred ranks per node and the value of the log posterior odds ratio per node as a function of the average degree of each network. To standardize the number of ranks per node within the interval $[0,1]$, we rescale it as $(R-1)/(N-1)$, where $N$ is the number of nodes in the network and $R$ is the total number of ranks inferred by the best-performing model. This quantity will be $1$ for perfect orderings in which $R=N$. It will be 00 if no rankings are inferred and all nodes are assigned the same strength, and it will take intermediate values when non-trivial partial rankings are present. The data reveals three distinct phases based on the density of the underlying network. For low average degree values, $\langle k\rangle$, the evidence is insufficient to support any meaningful ranking, leading to all actors being assigned the same strength. As connectivity increases, partial rankings begin to offer a more parsimonious description of the data compared to a perfect ordering of the nodes. In this intermediate regime, non-trivial subdivisions emerge, where actors are grouped by strength: all nodes within a group share the same strength, but strengths vary across groups. Finally, as the networks become sufficiently dense, enough data is available to support a perfect ordering of the nodes, and the BT model provides a more parsimonious description of the data.
@@
where $n_{r}$ is the size of rank $r$ and $N$ is the total number of nodes in the network. This measure reflects the balance in the size distribution of rankings: $R^{*}$ will achieve a maximum value of $R$ when all the rankings are of equal size. In contrast, it approaches a minimal value close to one when the distribution is highly imbalanced, with most nodes concentrated in a single rank. Fig. 3(b) shows the $R^{*}$ as a function of the inferred partial rankings $R$ for the PR model. While for most datasets $R^{*}\simeq R$, indicating that in most cases the PR algorithm infers roughly equally sized rankings, there are some notable exceptions, most notably the dataset of ATP tennis matches between 2010 and 2019, which displays a considerably smaller effective number of groups, reflecting a pronounced imbalance in the sizes of the inferred ranks. Specifically, the strongest 15 players are distributed across three partial rankings, while over $84\%$ of all players are grouped into the weakest rank. This skewed decomposition likely stems from the structure of professional tennis tournaments. ATP matches are generally organized in a knock-out format, where losers are progressively eliminated. While this structure is efficient and straightforward, it can result in noisy rankings, as strong players may be eliminated early due to random fluctuations or unfavorable matchups. To mitigate this, the ATP organizes multiple knock-out tournaments throughout the year, ensuring that the strongest players consistently rise to the top despite potential variability in individual tournaments. However, this system disproportionately concentrates matches among the strongest players, who predominantly compete against one another as they progress through the brackets. Consequently, the dataset contains ample information to distinguish the strongest players but far less data to differentiate among weaker players, leading to highly unbalanced partial rankings in terms of size.
@@
Finally, we also compare the results inferred by our algorithm with those obtained by the BT model. Fig. 3(d) shows the number of nodes $N$ in each network as a function of the rankings $R$ inferred by the partial rankings algorithm, where the points are colored according to the $\tau_{B}$ score between the rankings inferred by the PR and BT models. Again, we observe a general tendency for the $\tau_{B}$ scores to increase as $R$ approaches $N$.

### v3

@@ II.4 Optimization
[…]
@@
We propose a fast nonparametric agglomerative algorithm based on an alternating optimization strategy to approximate the MAP estimates $\bm{\hat{r}}$, $\bm{\hat{\sigma}}$ for the node rankings and strengths. Starting from an initial condition in which each node belongs to its own group, so that $R=N$, the algorithm alternates between updating the unique strengths $\bm{\sigma}$ and rankings $\bm{r}$ so as to minimize the negative log-posterior $\mathcal{L}$ (Eq. (II.3)).
@@
Eq. (28) defines a set of $R$ self-consistent equations for the $\sigma_{r}$’s which can be iteratively solved until convergence. If $E(\bm{r})$ is the number of non-zero entries in the matrix $\bm{\omega}$ when the ranks are equal to $\bm{r}$, then the complexity of this algorithm is $O(E(\bm{r}))$. These iterative updates for the player’s strengths are of the same form as those introduced in Eq. (12) and benefit from the same convergence guarantees.
@@
which is the standard iterative update of Eq. (28) and can be solved efficiently with a computational complexity of $O(E(\bm{r})/R)$, where $E(\bm{r})$ is now the number of non-zero entries in the matrix $\bm{\omega}$ for the current ranks $\bm{r}$. When applied across all R1R-1 possible adjacent rank pairs $(r,r^{\prime})$, the total complexity for this update and merge process is then $O(E(\bm{r}))$.
@@
Fig. 1 shows the results of applying both the BT model and our Partial Rankings (PR) algorithm (with the BT likelihood of Eq. (3)) to a small dataset of dominance interactions in a pack of wolves van198711 . We notice that our model offers a more concise interpretation of the data by grouping individuals with similar BT scores into the same rankings.
@@ III.1 Synthetic Match Datasets
Figure 2: (a) Heatmap of the number of rankings $R$ inferred by the partial rankings model, as a function of $(\sigma,\langle m\rangle)$ for a dataset of $N=50$ nodes planted into $R=3$ rankings as described by the synthetic model in Sec. III.1. (b) Heatmap of the log posterior odds ratio (Eq. (38)) between the BT and the partial rankings model across the simulations. Positive values indicate a preference for the partial rankings model, and negative values a preference for the BT model.
Figure 3: (a) Average number of rankings inferred by the best-performing model (blue) and our partial rankings model (red) as a function of the average number of matches per player pair $\langle m\rangle$ for the case in which no planted partial rankings are present. (b) Log-posterior-odds ratio (normalized per node) between the two models. Negative values of this difference indicate a preference for the BT model and positive values a preference for the partial rankings model. All results were obtained by averaging over 2020 different simulations from the synthetic network model of Sec. III.1, and error bars indicate 22 standard errors from the mean. Except near the transition point, error bars are smaller than the marker size.
@@
    For a given value of $\sigma$, assign three scores $\sigma,0,\sigma]$ (representing player strengths $e^{-\sigma},1,e^{\sigma}]$) to define the planted partial rankings.
@@
By varying $M$ and $\sigma$, we generate networks of matches with different levels of sparsity and strength separation, allowing us to assess the performance of our algorithm across a wide range of scenarios. While the networks produced by this synthetic model differ from those typically observed in real data—in terms of density, clustering, or other structural features—working with such simplified models allows us to isolate and systematically examine how specific factors, namely the number of matches per pair and the separation of underlying player strengths, influence the algorithm’s performance. To that end, we adopt the simplest possible null model, in which matches are generated as random graphs, leaving all other structural features unspecified.
[…]
$\langle m\rangle=\frac{M}{\binom{N}{2}}$, (37)
for $N=50$ and different total numbers of matches $M$.
For each $(\sigma,\langle m\rangle)$ pair, $20$ different networks were generated, and the results were then averaged. The ranges of values for $\langle m\rangle$ and $\sigma$ were chosen so as to capture all the relevant behaviors of the model.
As expected, we observe that for low values of $\langle m\rangle$, there is not enough evidence to recover the planted rankings regardless of the strength separation between players. In such sparse conditions, the algorithm assigns all players to the same rank, as shown in Fig. 2(a). This behavior contrasts sharply with the BT model, which still produces a complete ranking of the nodes even in these highly sparse conditions. As the number of matches increases and the networks become denser, the model gains sufficient statistical evidence to distinguish between the rankings and eventually recovers the planted structure. Notably, the number of matches required to correctly recover the rankings decreases rapidly with greater strength separation between players. In these cases, stronger players consistently outperform weaker ones, creating clearer signals that enable the ranking structure to be resolved with less data.
@@
as a function of $\langle m\rangle$ and $\sigma$. We observe that, when the separation in player strengths is small, our partial rankings model outperforms BT in sparse networks. In these cases, the PR model effectively leverages its ability to group players into shared ranks, avoiding overfitting. However, as the networks become denser, while our model accurately recovers the planted rankings even in this low-separation regime, it does not offer a more parsimonious description of the data (in terms of the posterior odds ratio of Eq. (38)) compared to the standard BT model which, in contrast, infers a fully ranked structure for the nodes with $R=N=50$. This suggests a noisy regime where the evidence supporting partial rankings is relatively weak. The BT likelihood provides a more flexible fit for fine-gained variations in the data in these dense regimes (effectively overfitting the data), even if its prior poorly reflects the underlying data.
@@
In Fig. 3 we analyze the performance of both models in the absence of partial rankings, using networks of $N=150$ players, each assigned a unique score sampled from a logistic distribution with mean zero and scale one—a setting in which the BT model is expected to accurately capture the data. Averaged over 20 network realizations, the results show the behaviour of our PR model and the best performing model, defined as the model (BT or PR) with the smallest negative log-posterior. For small values of $\langle m\rangle$, our PR model offers a more parsimonious fit, as the limited number of interactions prevents reliable inference of a complete ranking. Consequently, the log-posterior-odds ratio is positive in this regime, and the best performing model coincides with PR. As $\langle m\rangle$ increases, however, we observe a sharp transition in which the BT model quickly surpasses the PR model in explanatory power—the log-posterior odds ratio becomes negative, and the best performing model switches to BT, which infers $R=N=150$ distinct rankings. We also observe that the PR model continues to infer more rankings as $\langle m\rangle$ grows, albeit very slowly, indicating that, in principle, it is capable of recovering the full ranking given sufficient data. We note that care has to be taken with large values of $R$ at high $\langle m\rangle$, as, in these large data regimes the influence of the prior terms on the log posterior diminishes and so does the regularization they provide. See Appendix C for a more detailed analysis of the behavior of the model for large $R$ and $\langle m\rangle)$.
@@
Interestingly, as the inferred number of groups increases with $\langle m\rangle$, the log posterior odds ratio decreases. This is somewhat counterintuitive, as one might expect the posterior odds ratio to improve as the number of rankings inferred by the PR model approaches the true number of rankings. However, as mentioned in Sec. II, the PR model tends to heavily penalize the presence of a large number of rankings so that the net effect of increasing the number of rankings is a reduction in the posterior odds ratio, with the PR model becoming a less efficient encoding for the rankings as the number of inferred groups grows. In practice, the best practice is to evaluate both an original ranking method and our corresponding partial ranking method to determine which provides a more parsimonious fit to the data in terms of log-posterior odds.
@@
To assess the goodness of the recovered partitions, in Appendix D we compute the Kendall rank correlation coefficient ($\tau_{B}$), adjusted for ties, between the ground truth rankings and the rankings inferred by the BT and PR models. This comparison is performed both when partial rankings are present and when they are not. We observe that when partial rankings are present in the network, the PR model can quickly attain perfect recovery as the networks become denser. On the other hand, while the BT model still achieves high $\tau_{B}$ scores, it never achieves perfect ordinal association and always underperforms relative to the PR model. However, when no partial rankings are present, and each player is assigned a unique score, the roles are reversed, and BT always displays a higher ordinal association with the ground truth rankings. We also compare the player strengths inferred by the PR and BT models to the ground truth, finding that both models recover scores that closely match the planted values. However, the PR model is able to leverage the coarse grained structure of the player strengths to group them into partial rankings.
@@
We also compare our model with the partial rankings derived by applying 1D mean shift (MS) clustering comaniciu2002mean to the rankings inferred by the BT model in Appendix E. We observe that our algorithm can always recover the correct number of partial rankings when these are present, while MS fails to do so. More importantly, our algorithm can adjust the number of inferred clusters as more data becomes available, even in cases when no partial rankings exist, resulting in a more accurate description of the rankings. In contrast, mean shift clustering consistently infers roughly the same number of partial rankings regardless of the evidence provided by the data.
@@ III.2 Real Match Datasets
Figure 4: Partial rankings in real networks of pairwise comparisons. (a) (top) Rescaled number of ranks per node inferred by the best-performing model and (bottom) log posterior odds ratio per node, both as a function of $\langle m\rangle$ for all real-world networks considered in the study (see Table 1). The colours of the points indicate the different categories to which the datasets belong, while the shape of the markers indicates which model (PR or BT) emerged as the best-performing method (in terms of the posterior odds ratio). Positive values of the log posterior odds ratio indicate evidence in favor of the PR model; negative values indicate evidence in favor of the BT model. (b) Effective number of ranks, $R^{*}$ (Eq. (39)), as a function of the number of unique ranks $R$ inferred by the PR algorithm. The black dashed line represents the line $R^{*}=R$. (c) Number of rankings inferred via the partial rankings algorithm as a function of the number of rankings inferred via mean shift clustering. Point colors indicate the $\tau_{B}$ score between the two inferred rankings. The black dashed line represents the line $R_{PR}=R_{MS}$. (d) Number of nodes of each network as a function of the number of rankings $R$ inferred by the partial rankings algorithm. Point colors indicate the $\tau_{B}$ score between the rankings inferred via the PR algorithm and those inferred by the BT model. The dashed black line indicates the line $R=N$.
Data set NN $M$ $\langle m\rangle$ $R$ $R^{*}$ $\Delta\mathcal{L}$ Category Description Ref.
Soccer 2204 7438 0.003 1 1 1469.3 Sports/games Men’s international association football matches 2010 - 2019 soccer
Friends 774 2799 0.009 2 2.0 450.1 Human High-school friend nominations high_school_friendships
Tennis 1272 29397 0.036 6 1.7 404.9 Sports/games Association of Tennis Professionals matches 2010–2019 tennis
Chess 917 7007 0.017 1 1 357.5 Sports/games Online chess games on lichess.com in 2016 chess
@@
Little Rock food web 183 2494 0150 5 3.2 83.7 Animal Food web among the species in Little Rock Lake in Wisconsin martinez1991artifacts
@@
CS departments 205 34388 0.210 5 3.6 33.4 Human PhD graduates of one department hired as faculty in another clauset2015systematic
College football 115 593 0.090 1 1 26.8 Sports/games NCAA College Football matches 2013-2023 collegefootball
@@
Dutch school friends 26 234 0.720 2 1.3 9.2 Human Friendships at secondary school in The Netherlands, 2003-2004 snijders2010introduction
@@
Video games 125 1951 0.252 3 2.2 4.7 Sports/games Super Smash Bros Melee tournament matches in 2022 videogames
@@
History departments 144 4112 0.400 6 3.8 -3.2 Human PhD graduates of one department hired as faculty in another clauset2015systematic
@@
Hyenas 29 1913 4.712 9 7.9 -7.6 Animal Dominance interactions among hyenas in captivity strauss2019social
@@
Sparrows 26 1238 3.809 8 7.3 -15.4 Animal Attacks and avoidances among sparrows in captivity watt1986relationship
@@
Baboons 53 4464 3.239 13 10.3 -16.3 Animal Dominance interactions among baboons in captivity franz2015self
@@
Dogs 27 1143 3.256 6 5.3 -20.3 Animal Aggressive behaviors in a group of domestic dogs silk2019elevated
Wolf 15 10382 98.876 10 8.6 -24.0 Animal Dominance interaction among wolves in captivity van198711
@@
Mice 30 1230 2.828 5 4.2 -26.8 Animal Dominance interactions among mice in captivity williamson2016mouse
@@
Business departments 112 7856 1.264 9 7.3 -35.2 Human PhD graduates of one department hired as faculty in another clauset2015systematic
@@
Vervet monkeys 41 2980 3.634 8 6.8 -42.7 Animal Dominance interactions among a group of wild vervet monkeys vilette2020comparing
Scrabble 587 23477 0.137 2 2.0 -86.0 Sports/games Scrabble tournament matches 2004–2008 scrabble
Basketball 240 10002 0.349 2 2.0 -90.7 Sports/games National Basketball Association games 2015–2022 basketball
Soccer (aggregated) 288 7438 0.180 6 4.8 -131.1 Sports/games Soccer dataset above aggregated across 2010 - 2019 soccer
@@
Table 1: Datasets analysed using the PR and BT algorithms, in order of decreasing log posterior odds ratio in Eq. (38) (i.e. increasingly in favor of the original BT model). NN, $M$, and $\langle m\rangle$ indicate the number of nodes, edges, and density of each network, respectively. $R$ and $R^{*}$ indicate the number of partial rankings inferred via the PR algorithm and the effective number of partial rankings (Eq. (39)). $\mathcal{L}$ indicates the logarithm of the posterior odds ratio between the PR and BT models (Eq. (38)).
@@
In this section, we focus on how our PR algorithm may be used to derive patterns in real-world data. We consider the datasets compiled in jerdee2024luck and available at jerdeegit . In addition, we also consider a dataset of social hierarchies among a pack of wolves van198711 ; newman2023efficient , and other sets of directed networks available at netzschleuder . The datasets are summarized in Table 1.
@@
We apply our PR algorithm and the BT model to all match lists in the dataset and compare the results. In Fig. 4(a), we show the number of inferred ranks per node and the value of the log posterior odds ratio per node as a function of $\langle m\rangle$ for each network. To standardize the number of ranks per node within the interval $[0,1]$, we rescale it as (R1)/(N1)(R-1)/(N-1), where $N$ is the number of nodes in the network and $R$ is the total number of ranks inferred by the best-performing model. This quantity will be $1$ for perfect orderings in which $R=N$. It will be 0 if no rankings are inferred and all nodes are assigned the same strength, and it will take intermediate values when non-trivial partial rankings are present. The data reveals three distinct phases based on the density of the underlying network. For low densities, $\langle m\rangle$, the evidence is insufficient to support any meaningful ranking, leading to all actors being assigned the same strength. As connectivity increases, we observe an intermediate phase, in which partial rankings can often offer a more parsimonious description of the data compared to a perfect ordering of the nodes. In this intermediate regime, non-trivial subdivisions emerge, where actors are grouped by strength: all nodes within a group share the same strength, but strengths vary across groups. Finally, as the networks become sufficiently dense, enough data is available to support a perfect ordering of the nodes, and the BT model provides a more parsimonious description of the data.
@@
where $n_{r}$ is the size of rank $r$ and $N$ is the total number of nodes in the network. This measure reflects the balance in the size distribution of rankings: $R^{*}$ will achieve a maximum value of $R$ when all the rankings are of equal size. In contrast, it approaches a minimal value close to one when the distribution is highly imbalanced, with most nodes concentrated in a single rank. Fig. 4(b) shows $R^{*}$ as a function of the inferred partial rankings $R$ for the PR model. While for most datasets $R^{*}\simeq R$, indicating that in most cases the PR algorithm infers roughly equally sized rankings, there are some notable exceptions, most notably the dataset of ATP tennis matches between 2010 and 2019, which displays a considerably smaller effective number of groups, reflecting a pronounced imbalance in the sizes of the inferred ranks. Specifically, the strongest 15 players are distributed across three partial rankings, while over $84\%$ of all players are grouped into the weakest rank. This skewed decomposition likely stems from the structure of professional tennis tournaments. ATP matches are generally organized in a knock-out format, where losers are progressively eliminated. While this structure is efficient and straightforward, it can result in noisy rankings, as strong players may be eliminated early due to random fluctuations or unfavorable matchups. To mitigate this, the ATP organizes multiple knock-out tournaments throughout the year, ensuring that the strongest players consistently rise to the top despite potential variability in individual tournaments. However, this system disproportionately concentrates matches among the strongest players, who predominantly compete against one another as they progress through the brackets. Consequently, the dataset contains ample information to distinguish the strongest players but far less data to differentiate among weaker players, leading to highly unbalanced partial rankings in terms of size.
@@
Finally, we also compare the results inferred by our algorithm with those obtained by the BT model. Fig. 4(d) shows the number of nodes $N$ in each network as a function of the rankings $R$ inferred by the partial rankings algorithm, where the points are colored according to the $\tau_{B}$ score between the rankings inferred by the PR and BT models. Again, we observe a general tendency for the $\tau_{B}$ scores to increase as $R$ increases.
Across real datasets from various domains we have found the emergence of three distinct regimes in rank detectability as a function of the network density. For low densities, there is too little statistical evidence available for distinguishing rankings and we infer a single partial rank, with the partial rankings model being preferred over the standard BT model. In an intermediate density regime of around $\langle m\rangle\simeq 0.02$ matches per node pair, we find a modest number of partial ranks much less than $N$, and that the partial rankings model is often more parsimonious than BT in terms of posterior odds. Finally, for higher densities, we find that the BT model is most parsimonious, so that partial rankings are no longer needed as there is sufficient evidence in the data from the matches to distinguish the rankings of the nodes. These results are consistent with those in Sec. III.1. We also find of interest the considerable heterogeneity in inferred cluster sizes, suggesting that the underlying topology of the interaction network can influence the inferred partial rankings. This is of particular interest for ranking models such as BT that do not model edge placement, i.e. how the matches are decided. We additionally identify substantial discrepancies between the results of our pairwise partial ranking and those obtained by simply performing 1D clustering of the ranks, which highlights the importance of properly regularized ranking approaches that infer ties simultaneously while inferring the rankings. These partial rankings are shown to have increased correlation with the standard BT ranks as the number of ties approaches zero.